# On the Welfare of EIP-1559 with Patient Bidders

### v1

@@ 5 Model Extensions
In this section we consider two model extensions: to partially patient bidders and to multiple resources. In both cases we present impossibilities, showing that losing a constant fraction of the welfare in inevitable even for general online algorithms (not only price-based algorithms or the EIP-1559 algorithm).
@@ 5.1 Partially Patient Bidders: Online Algorithms must Lose a Constant Factor
A transaction with value $v$ and discount factor $\rho$ (where $10\leq\rho<$ 1) has value for its user of $v\cdot(1-\rho)^{t_{e}-t_{a}}$ if where $t_{a}$ is its arrival time and $t_{e}$ its execution time.
@@
A transaction with value $v$ and patience $p$ has value of $v$ for its user if it have been scheduled within $p$ blocks of its arrival time and value 00 otherwise. I.e. if $t_{a}$ is its arrival time and $t_{e}$ its execution time then the user’s value is $v$ if $t_{a}\leq t_{e}\leq t_{a}+p$ (and 00 otherwise).
@@
Fix an online algorithm with average block size limit $B$, slackness $\Delta$, and extension $\Gamma$, where $\Delta(T)+\Gamma(T)=o(T)$. For every minimum discount rate $\rho_{min}>0$ there exist a time horizon $T$, and an input sequence where all bidders have discount rates $\rho_{i}\leq\rho_{min}$ and values in $\{1,2\}$ for which the algorithm loses at least a fraction $\delta_{0}=1/20$ of welfare relative to the optimal schedule with worst-case block size limit $B$.
@@
Fix an online algorithm with average block size limit $1$, slackness $\Delta(T)$, and extension $\Gamma(T)$. Choose $T$ that is larger than $\Gamma(T)+\Delta(T)$ by a large factor $K$, and also such that $(1-\rho_{min})^{T/3}\leq 1/2$. Let $p=T/3$ and denote $\rho=\rho_{min}$. Now consider the following input scenario: At block 1, We have $p$ “patient” transactions arriving, each with size $q_{i}=1$, value $v_{i}=2$, and discount rate $\rho_{i}=0$. Also at each of the first $p$ blocks arrives one “hasty” transactions, each with size $q_{i}=1$, value $v_{i}=1$, and discount rate $\rho_{i}=\rho$. So now our algorithm has to decide what to schedule during the first $p$ steps: the high-value patient transactions or the low-value hasty transactions.
@@
Case I: by the end of time $p$ at least $p/2$ hasty transactions have been scheduled. In this case, for block $p+1$ our adversary will choose as input $2p$ more transactions each with size $q_{i}=1$, value $v_{i}=2$, and discount rate $\rho_{i}=0$. Now by the end of time $T=3p$ at most $T+\Delta(T)+\Gamma(T)=T\cdot(1+O(1/K))$ transactions have been scheduled, at least $p/2$ of them of value $1$ and the rest of value $2$. So the total social welfare achieved is at most $2\cdot(5/2+O(1/K))p+1\cdot p/2=(11/2+O(1/K)p$. On the other hand, the optimal offline benchmark would only schedule the high-value patient transactions for a social welfare of $2\cdot 3p=6p$, so the ratio of welfares is $11/12+O(1/K)$ which, by choosing $K$ large enough, is less than $19/20$.
@@
Case II: on the end of time $p$ at most $p/2$ hasty transactions have been scheduled. In this case, in each of the next $p$ steps we get a single transaction with size $q_{i}=1$, value $v_{i}=1$, and discount rate $\rho_{i}=\rho$ (and nothing more arrives in the last $p$ steps). We can assume wlog that our online algorithm was “clever enough” to schedule all new transactions during the “middle” $p$ steps as otherwise switching a scheduled hasty transaction or patient transaction with the new transaction that arrived at that time step will only increase the social welfare (since both the hasty transaction and the new transaction lose value at the same rate $\rho$ but the new transaction has not lost value yet when it arrived while the hasty transactions have already lost some of their initial value. Patient transactions do not lose value at all.) It follows that at most $\Delta$ hasty transactions where scheduled during the ”middle” $p$ blocks and so the social welfare that our algorithm gets can be bounded from above by $(19/4+o(1))\cdot p$: $2p$ from all the new transactions (that are scheduled in the middle $p$ steps) plus $2p$ from all the patient transactions plus at most $3p/4+\Delta$ from the hasty transactions since at least half of them (minus $\Delta$) are scheduled in the last $p$ block after they have lost at least half their value (since each of them waited at least for the middle $p$ steps and $(1-\rho)^{p}\leq 1/2$). The optimal offline solution is to schedule the hasty transactions as they come in the first $p$ steps, schedule the new transactions, as they come, in the middle $p$ steps, and leave the patient transactions for the last $p$ steps, getting the full value of $5p$. The ratio between these welfares is at most $19/20+o(1)$. ∎
@@
Notice that the proof uses transactions with different discount rates. It is interesting to ask what can be done when all transactions share the same, global, discount rate. We conjecture that, in fact, the modified EIP-1559 (that takes into account current values) does produce nearly-optimal schedules with a loss that approaches the loss of the fully patient case, as the minimum discount rate $\rho_{min}$ approaches 00.
@@
Fix an online algorithm with average block size limit $B$, slackness $\Delta$, and extension $\Gamma$. For every global patience level $p$ and every time horizon $T\geq 2p$, there exists an input sequence where all bidders have patience $p_{i}=p$ and values in $\{1,2\}$ where the algorithm loses at least a fraction $\delta_{0}=1/10$ of welfare relative to the optimal schedule with worst-case block size limit $B$.
@@
Let $B=1$ and fix an online algorithm with average block size limit $B$ and slackness $\Delta$. Let $p=(\Gamma+\Delta)/\varepsilon$ for some small constant $\varepsilon$ to be specified below. We will prove the lemma for $T=2p$ and the same follows for any larger $T$ by adding an arbitrarily long time prefix with no input transactions. At time 1 arrive $p$ “green” transactions with value 1 and size 1. Additionally, every time step for the next $p-1$ steps arrives a single“red” transaction with value 2 and size 1. Let us look at how many red transactions were scheduled by our algorithm by the end of time $p$.
Case I: at least $p/2$ red transactions were scheduled. Since the total number of scheduled transactions by this time is at most $p+\Delta$, this means that at least $p/2-\Delta$ green transactions were not scheduled and can never be scheduled any more since their deadline has passed. Now assume that no more transactions arrive in the next $p$ steps, so our algorithm can schedule the rest of the red transactions but, again, not the unscheduled green ones. This gives an upper bound of $2p+p/2+\Delta$ on the social welfare of our online algorithm. The Optimal clairvoyant algorithm could have scheduled only the green transactions during the first $p$ steps and then the red transactions, one after another, after that for a total social welfare of $2(p-1)+p$. The ratio between these amounts is $5/6+O(\varepsilon)$.
Case II: less than $p/2$ red transactions were scheduled by the end of time $p$. Now assume that another $p$ “new” transactions arrive at time $p+1$ each with value 2 and size 1. Since the total quantity scheduled in the first $p$ steps is at most $p+\Delta$ and since at most $p/2-1$ of this quantity has a value per unit of 2 and the rest have value 1, then during the first $p$ steps our online algorithm could have made at most $p+\Delta+p/2-1$ social welfare. In the second half, including the extra allowed $\Gamma$ steps, an upper bound to the social welfare that it can make is $2(p+\Delta+\Gamma)$. Thus the total welfare of our online algorithm is bounded from above by $7p/2+O(\varepsilon)$ welfare. The Optimal clairvoyant algorithm could have scheduled only the red transactions during the first $p$ steps and then scheduled the new transactions in the next $p$ steps for a total welfare of $4p$. The ratio between these amounts is $7/8+O(\varepsilon)$. Choosing $\varepsilon$ small enough that this amount is less than $1/10$ concludes the proof of the theorem. ∎
@@ 5.2 Multidimensional Case: Online Algorithms Must Lose a Constant Factor
Consider the case of three resources, X, Y, and Z, and assume that the block size for each resource is $B_{X}=B_{Y}=B_{Z}=1$. Let $t=T/2$ where $T$ is the horizon over which our online algorithm aims to approximate OPT and fix at an online algorithm that uses each resource at total capacity of at most $(t+\Delta)$ during any consecutive $t$ blocks.
@@
During the first $t$ steps there is only capacity of $t+\Delta$ of resource $Z$ so at least one of the two types of demanded bundles is allocated at most at total capacity $(t+\Delta)/2$. Without loss of generality assume that the bundle $\{X,Z\}$ is allocated at most $(t+\Delta)/2$ during the first $t$ steps. Now at time $t+1$ we get another demand of $t$ for the singleton bundle $\{X\}$ again at value 1 per unit of $X$. (Of course, had the bundle $\{Y,Z\}$ been the one that was at least half-un-allocated then our adversary would have brought demand for the singleton bundle $\{Y\}$ instead.) The best that our online algorithm can do at this point is pack singleton bundles $X$ together with any leftover of the original bundles $\{Y,Z\}$ into the remaining blocks, and leaving at least $t/2-\Delta-\Gamma$ demand for the bundle $\{X,Z\}$ unsatisfied for a total value of at most $(5t+\Delta+\Gamma)/2$. OPT, on the other hand, could have fully allocated everything by allocating only the $\{X,Z\}$ bundles for the first $t$ blocks and packing the demand for $\{Y,Z\}$ with $\{X\}$ in each of the last $t$ blocks getting a total value of $3tB$. The ratio between these two levels of social welfare is $5/6+o(1)$. ∎
We do not know whether the case of two resources has a similar impossibility result, we leave this problem, and the problem of getting some constant-factor approximation with any number of resources, for future work.
@@ Acknowledgments
We thank Shiri Ron for useful discussions during the early stages of this work. Noam Nisan thanks his colleagues in Starkware for many useful discussions. Moshe Babaioff’s research is supported in part by a Golda Meir Fellowship and the Israel Science Foundation (grant No. 301/24). Noam Nisan’s research was supported by a grant from the Israeli Science Foundation (ISF number 505/23).

### v2

@@ 5 Model Extensions
In this section we consider two model extensions: to partially patient bidders and to multiple resources. In both cases we present impossibilities, showing that losing a constant fraction of the welfare is inevitable even for general online algorithms (not only price-based algorithms or the EIP-1559 algorithm).
@@ 5.1 Partially Patient Bidders: Online Algorithms must Lose a Constant Factor
A transaction with value $v$ and discount factor $\rho$ (where $0\leq\rho<1$) has value for its user of $v\cdot(1-\rho)^{t_{e}-t_{a}}$, where $t_{a}$ is its arrival time and $t_{e}$ its execution time.
@@
A transaction with value $v$ and patience $p$ has value of $v$ for its user if it has been scheduled within $p$ blocks of its arrival time and value 0 otherwise. I.e. if $t_{a}$ is its arrival time and $t_{e}$ its execution time then the user’s value is $v$ if $t_{a}\leq t_{e}\leq t_{a}+p$ (and 0 otherwise).
@@
Fix an online algorithm with average block size limit $B$, slackness $\Delta$, and extension $\Gamma$, where $\Delta(T)+\Gamma(T)=o(T)$. For every maximum discount rate ${\rho_{max}}>0$ there exist a time horizon $T$, and an input sequence where every bidder $i$ has a non-negative discount rate $\rho_{i}\in\{0,\rho_{max}\}$ and value $v_{i}\in\{1,2\}$, for which the algorithm loses at least a fraction $\delta_{0}=1/20-o(1)$ of welfare relative to the optimal schedule with worst-case block size limit $B$.
@@
Fix an online algorithm with average block size limit $1$, slackness $\Delta(T)$, and extension $\Gamma(T)$. Choose $T$ that is larger than $\Gamma(T)+\Delta(T)$ by a large factor $K$, and also such that $(1-{\rho_{max}})^{T/3}\leq 1/2$. Let $p=T/3$ and denote $\rho={\rho_{max}}$. Now consider the following input scenario: At block 1, we have $p$ “patient” transactions arriving, each with size $q_{i}=1$, value $v_{i}=2$, and discount rate $\rho_{i}=0$. Also at each of the first $p$ blocks arrives one “hasty” transaction, each with size $q_{i}=1$, value $v_{i}=1$, and discount rate $\rho_{i}=\rho$. So now our algorithm has to decide what to schedule during the first $p$ steps: the high-value patient transactions or the low-value hasty transactions.
@@
Case I: by the end of time $p$ at least $p/2$ hasty transactions have been scheduled. In this case, for block $p+1$ our adversary will choose as input $2p$ more transactions each with size $q_{i}=1$, value $v_{i}=2$, and discount rate $\rho_{i}=0$. Now by the end of time $T=3p$ at most $T+\Delta(T)+\Gamma(T)=T\cdot(1+O(1/K))$ transactions have been scheduled, at least $p/2$ of them of value $1$ and the rest of value $2$. So the total social welfare achieved is at most $2\cdot(5/2+O(1/K))p+1\cdot p/2=(11/2+O(1/K))p$. On the other hand, the optimal offline benchmark would only schedule the high-value patient transactions for a social welfare of $2\cdot 3p=6p$, so the ratio of the two welfare terms is $11/12+O(1/K)$ which, by choosing $K$ large enough, is less than $19/20$.
@@
Case II: by the end of time $p$ at most $p/2$ hasty transactions have been scheduled. In this case, in each of the next $p$ steps we get a single transaction with size $q_{i}=1$, value $v_{i}=2$, and discount rate $\rho_{i}=\rho$ (and nothing more arrives in the last $p$ steps). We can assume wlog that our online algorithm was “clever enough” to schedule all new transactions during the “middle” $p$ steps as otherwise switching a scheduled hasty transaction or patient transaction with the new transaction that arrived at that time step will only increase the social welfare (since both the hasty transaction and the new transaction lose value at the same rate $\rho$ but the new transaction has higher current value. Patient transactions do not lose value at all.) It follows that at most $\Delta$ hasty transactions were scheduled during the “middle” $p$ blocks and so the social welfare that our algorithm gets can be bounded from above by $(19/4+o(1))\cdot p$: $2p$ from all the new transactions (that are scheduled in the middle $p$ steps) plus $2p$ from all the patient transactions plus at most $3p/4+\Delta$ from the hasty transactions since at least half of them (minus $\Delta$) are scheduled in the last $p$ blocks after they have lost at least half their value (since each of them waited at least for the middle $p$ steps and $(1-\rho)^{p}\leq 1/2$). The optimal offline solution is to schedule the hasty transactions as they come in the first $p$ steps, schedule the new transactions, as they come, in the middle $p$ steps, and leave the patient transactions for the last $p$ steps, getting the full value of $5p$. The ratio between these welfares is at most $19/20+o(1)$, and thus the loss is at least 1/20o(1)1/20-o(1). ∎
@@
Notice that the proof uses transactions with different discount rates. It is interesting to ask what can be done when all transactions share the same, global, discount rate. We conjecture that, in fact, the modified EIP-1559 (that takes into account current values) does produce nearly-optimal schedules with a loss that approaches the loss of the fully patient case, as the maximum discount rate $\rho_{max}$} approaches 0.
@@
Fix an online algorithm with average block size limit $B$, slackness $\Delta(T)=o(T)$, and extension $\Gamma(T)=o(T)$ then there exists $p^{*}$ such that for every global patience level $p\geq p^{*}$ and every sufficiently large time horizon $T\geq 2(p+1)$, there exists an input sequence where every bidder $i$ has patience $p_{i}=p$ and value $v_{i}\in\{1,2\}$, for which the algorithm loses at least a fraction $\delta_{0}=1/10$ of welfare relative to the optimal schedule with worst-case block size limit $B$.
@@
Without loss of generality assume that $B=1$ and fix an online algorithm with average block size limit $B$ and slackness $\Delta(T)$ and extension $\Gamma(T)$. Let $p^{*}$ be such that $\Delta(T)+\Gamma(T)\leq\varepsilon\cdot T$ for all $T\geq p^{*}$ for some small $\varepsilon$ to be chosen below. Given $p\geq p*$, we will prove the claim for $T=2(p+1)$ and the same applies for any larger $T$ by adding an arbitrarily long time prefix with no input transactions.
At time 1 arrive $p+1$ “green” transactions with value 1 and size 1. Additionally, every time step for the next $p$ steps arrives a single “red” transaction with value 2 and size 1. Let us look at how many red transactions were scheduled by our algorithm by the end of time $p+1$.
Case I: at least $(p+1)/2$ red transactions were scheduled. Since the total number of scheduled transactions by this time is at most $(p+1)+\Delta(T)$, this means that at least $(p+1)/2-\Delta(T)$ green transactions were not scheduled and can never be scheduled any more since their deadline has passed. Now assume that no more transactions arrive in the next $p+1$ steps, so our algorithm can schedule the rest of the red transactions but, again, not the unscheduled green ones. This gives an upper bound of $2(p+1)+(p+1)/2+\Delta(T)=5(p+1)/2+\Delta(T)$ on the social welfare of our online algorithm. The optimal offline algorithm could have scheduled only the green transactions during the first $p+1$ steps and then the red transactions, one after another, after that for a total social welfare of 2p+(p+1)=3p+1{2p+(p+1)=3p+1}. The ratio between these amounts is $5/6+O(\varepsilon)$, and thus the loss is at least 1/6o(1)>1/101/6-o(1)>1/10.
Case II: less than $(p+1)/2$ red transactions were scheduled by the end of time $p+1$. Now assume that another $p+1$ “new” transactions arrive at time p+2{p+2} each with value 2 and size 1. Since the total quantity scheduled in the first $p+1$ steps is at most $(p+1)+\Delta(T)$ and since at most (p+1)/21(p+1)/2-1 of this quantity has a value per unit of 2 and the rest have value 1, then during the first $(p+1)$ steps our online algorithm could have made at most $(p+1)+\Delta(T)+(p+1)/2-1$ social welfare. In the second half, including the extra allowed $\Gamma(T)$ steps, an upper bound to the social welfare that it can make is $2((p+1)+\Delta(T)+\Gamma(T))$. Thus the total welfare of our online algorithm is bounded from above by $7(p+1)/2+O(\varepsilon\cdot p)$ welfare. The optimal offline algorithm could have scheduled a green transaction in round 1, then the $p$ red transactions during the next $p$ steps and then scheduled the new transactions in the next $(p+1)$ steps for a total welfare of $1+2p+2(p+1)=4p+3$. The ratio between these amounts is $7/8+O(\varepsilon)$. Choosing $\varepsilon$ small enough that this amount is less than $9/10$ (and thus the loss is at least $1/10$) concludes the proof of the theorem. ∎
@@ 5.2 Multidimensional Case: Online Algorithms Must Lose a Constant Factor
Consider the case of three resources, X, Y, and Z, and assume that the block size for each resource is $B_{X}=B_{Y}=B_{Z}=1$. Let $t=T/2$ where $T$ is the horizon over which our online algorithm aims to approximate OPT and fix an online algorithm that uses each resource at total capacity of at most $(t+\Delta)$ during any consecutive $t$ blocks.
@@
During the first $t$ steps there is only capacity of $t+\Delta$ of resource $Z$ so at least one of the two types of demanded bundles is allocated at most at total capacity $(t+\Delta)/2$. Without loss of generality assume that the bundle $\{X,Z\}$ is allocated at most $(t+\Delta)/2$ during the first $t$ steps. Now at time $t+1$ we get another demand of $t$ for the singleton bundle $\{X\}$ again at value 1 per unit of $X$. (Of course, had the bundle $\{Y,Z\}$ been the one that was at least half-un-allocated then our adversary would have brought demand for the singleton bundle $\{Y\}$ instead.) The best that our online algorithm can do at this point is pack singleton bundles $X$ together with any leftover of the original bundles $\{Y,Z\}$ into the remaining blocks, and leaving at least $t/2-\Delta-\Gamma$ demand for the bundle $\{X,Z\}$ unsatisfied for a total value of at most $(5t+\Delta+\Gamma)/2$. OPT, on the other hand, could have fully allocated everything by allocating only the $\{X,Z\}$ bundles for the first $t$ blocks and packing the demand for $\{Y,Z\}$ with $\{X\}$ in each of the last $t$ blocks getting a total value of $3t$. The ratio between these two levels of social welfare is $5/6+o(1)$. ∎
@@ Acknowledgments
We thank Shiri Ron for useful discussions during the early stages of this work. Noam Nisan thanks his colleagues in Starkware for many useful discussions. Moshe Babaioff’s research is supported by a Golda Meir Fellowship and the Israel Science Foundation (grant No. 301/24). Noam Nisan’s research was supported by a grant from the Israeli Science Foundation (ISF number 505/23).